# No License

### v3

@@ 4.2 GAT
$e_{u,v}=\bm{a}^{\top}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (15)
with the message from node $v$ to node $u$ proportional to $\exp\left(e_{u,v}\right)\cdot\bm{W}\bm{h_{v}}$. While the model is anisotropic in nature, messages are nevertheless only linearly transformed with node $u$ only determining the degree of contribution through the scalar $e_{u,v}$. Meanwhile, SIR-GCN directly works with the unnormalized attention mechanism in Eqn. 15 and allows the features of the query node to dynamically transform messages. Specifically, if $\sigma=\textsc{LeakyReLU}$ and $\bm{W_{R}}=\bm{A}$, one obtains
$\bm{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\bm{A}~{}\textsc{LeakyReLU}\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right)$ (16)
which shows Eqn. 15 becoming a contextualized message in the SIR-GCN model. Nevertheless, GAT and GATv2 may be recovered, up to a normalizing constant, with an appropriate choice of $\sigma$.
@@ 4.3 GIN
Likewise, within the proposed SIR-GCN model, one may explicitly add a residual connection in the combination strategy to obtain
$\bm{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\bm{h_{u}})+\sum_{v\in\mathcal{N}(u)}\bm{W_{R}}~{}\sigma\left(\bm{W_{Q}}\bm{h_{u}}+\bm{W_{K}}\bm{h_{v}}\right),$ (17)
where $\textsc{MLP}_{\text{Res}}$ is a learnable residual network. If $\textsc{MLP}_{\text{Res}}(\bm{h})=(1+\epsilon)\cdot\bm{h}$, $\sigma=\textsc{PReLU}(\alpha=1)$, $\bm{W_{Q}}=\bm{0}$, and $\bm{W_{R}W_{K}}=\bm{I}$, then
$\bm{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\bm{h_{u}}+\sum_{v\in\mathcal{N}(u)}\bm{h_{v}}$ (18)
is equivalent to GIN. Hence, SIR-GCN with residual connection encompasses GIN.
@@ 4.4 PNA
Furthermore, while SIR-GCN and PNA approach the problem of uncountable node features differently, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes. The key difference lies with PNA using a static (Brody et al., 2021) or linear message function $m$ which translates to
$m\left(\bm{h_{v}},\bm{h_{u}}\right)=\bm{W_{K}}\bm{h_{v}}+\bm{W_{Q}}\bm{h_{u}}=\bm{W_{K}}\bm{h_{v}}+\bm{b_{u}}.$ (19)
As a result, the influence of the query node on the aggregated neighborhood features is limited. For instance, when using mean, max, or min aggregators, the influence of the query node $u$ is restricted to the bias term $\bm{b_{u}}$. Moreover, with normalized moment aggregators, the bias term is effectively canceled out during the normalization process, further reducing the influence of the query node. Hence, PNA does not fully leverage its anisotropic nature, limiting its expressivity. In contrast, the dynamic nature of SIR-GCN allows for the non-linear embedding of the features of the query node $\bm{h_{u}}$ within the aggregated neighborhood features, thereby fully leveraging its anisotropic nature.
@@
where the modification lies in concatenating the label of the center node with every element of the multiset before hashing. This modification, while negligible when $\mathcal{H}$ is countable, becomes significant when $\mathcal{H}$ is uncountable as noted in the previous section. Thus, SIR-GCN inherits the theoretical capabilities (and limitations) of the 1-WL test.
@@ 4.6 SIR-GCN
Overall, SIR-GCN offers flexibility in two key dimensions of GNNs: aggregation strategy and message transformation. Consequently, it generalizes four prominent GNNs in literature—GCN, GraphSAGE, GAT, and GIN—ensuring that it is at least as expressive as these models. Notably, SIR-GCN sets itself apart from other GNNs as the first MPNN instance to incorporate both anisotropic and dynamic (i.e., contextualized) messages within the MPNN framework, making it well-suited for heterophilous tasks (Bronstein et al., 2021) while remaining adaptable to homophilous tasks.
@@
Figure 3: SIR-GCN encompasses GCN, GraphSAGE, GAT, and GIN.
In addition, SIR-GCN distinguishes itself from PNA by employing only a single aggregator that theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity. Nevertheless, its expressivity is maintained through contextualized messages, allowing it to inherit the representational capability of the 1-WL test.
@@ 5 Experiments
Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are conducted to highlight the expressivity of SIR-GCN. To ensure fair evaluation, models not employing complex architectural design or manually crafted features using domain knowledge are used as primary comparisons.
@@ 5.1 Synthetic datasets
DictionaryLookup (Brody et al., 2021) consists of bipartite graphs with $2n$ $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute. The task is to predict the value of query nodes by matching their attribute with the key nodes.
@@
Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $n$. SIR-GCN and GATv2 achieve perfect accuracy attributed to their anisotropic and dynamic nature. However, it is observed that GATv2 suffers from performance degradation in some trials. Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic and/or static nature. The results underscore the utility of a dynamic attentional or relational mechanism in capturing the relationship between the query and key nodes.
GraphHeterophily
GraphHeterophily is an original synthetic dataset. It consists of random directed graphs with each node assigned one of $c$ classes. The task is to count the number of directed edges connecting two nodes with distinct class labels within each graph.
Table 2: Test mean squared error on GraphHeterophily.
@@
Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $c$. SIR-GCN achieves an MSE loss of nearly zero attributed to its anisotropic and dynamic nature and sum aggregation. In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al. (2018a). Meanwhile, GIN and PNA successfully retain the graph structure but fail to learn the relationship between the labels of the query node and key nodes due to their static nature. The results illustrate the utility of anisotropic and dynamic models using sum aggregation even with countable node features.
@@ 5.2 Benchmark datasets
Benchmarking GNNs (Dwivedi et al., 2023) is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks. In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks. Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features. The performance metric for ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy. Dwivedi et al. (2023) provides more information regarding the individual datasets.
@@
Table 3 presents the mean and standard deviation of the test performance for SIR-GCN and comparable GNN models across the six benchmarks where the experimental set-up follows that of Dwivedi et al. (2023) to ensure fair evaluation. The results show that SIR-GCN consistently outperforms popular GNNs in literature. Notably, SIR-GCN also outperforms both PNA (Corso et al., 2020) and efficient graph convolution (EGC-M) (Tailor et al., 2021) which use multiple aggregators. This highlights the significance of contextualized messages in enhancing the expressivity of GNNs, complementing the discussion in the previous section.
ogbn-arxiv
ogbn-arxiv (Hu et al., 2020) is a benchmark dataset representing the citation network between all Computer Science (CS) arXiv papers indexed by Microsoft academic graph (Wang et al., 2020). Each node represents an arXiv paper and a directed edge represents a citation. The task is to classify each paper, based on its title and abstract, into the 40 subject areas of arXiv CS papers.
Table 4: Test accuracy on ogbn-arxiv.
Model GIANT-XRT (Chien et al., 2021) BoT (Wang et al., 2021) C&S (Huang et al., 2020) Others Accuracy Parameters
GATv2 0.7415 $\pm$ 0.0005 207,520
GraphSAGE 0.7435 $\pm$ 0.0014 546,344
SIR-GCN 0.7525 $\pm$ 0.0009 667,176
0.7574 $\pm$ 0.0020 697,896
LGGNN 0.7570 $\pm$ 0.0018 1,161,640
RevGAT
KD, DCN
0.7636 $\pm$ 0.0013 1,304,912
AGDN
self-KD
0.7637 $\pm$ 0.0011 1,309,760
Table 4 presents the mean and standard deviation of the test accuracy for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that SIR-GCN, utilizing only a single GNN layer, outperforms comparable models in predicting the subject area of the papers. As expected, however, SIR-GCN fails to compete with complex frameworks utilizing more tricks such as the reversible GAT (RevGAT) (Li et al., 2021) and the adaptive graph diffusion network (AGDN) (Sun et al., 2020), both of which build upon GAT by employing grouped reversible residual connections and adaptive graph diffusion, respectively. Nevertheless, SIR-GCN achieves performance close to that of the complex GNN frameworks mentioned, showcasing an effective balance between complexity and expressivity.
ogbg-molhiv
ogbg-molhiv (Hu et al., 2020) is another benchmark dataset where each graph represents a molecule with nodes representing atoms and edges representing chemical bonds. Node features contain information regarding the atom while edge features contain information regarding the chemical bond. The task is to predict whether or not the molecules inhibit HIV replication.
Table 5: Test ROC-AUC on ogbg-molhiv.
Model GraphNorm (Cai et al., 2021) VirtualNode (Gilmer et al., 2017) Others ROC-AUC Parameters
GIN
FLAG
0.7748 $\pm$ 0.0096 3,336,306
GIN 0.7773 $\pm$ 0.0129 1,518,901
EGC-M 0.7818 $\pm$ 0.0153 317,265
GCN 0.7883 $\pm$ 0.0100 526,201
PNA 0.7905 $\pm$ 0.0132 326,081
SIR-GCN 0.7721 $\pm$ 0.0110 327,901
0.7981 $\pm$ 0.0062 328,201
GSN 0.7799 $\pm$ 0.0100 3,338,701
GSAT 0.8067 $\pm$ 0.0950 249,602
CIN 0.8094 $\pm$ 0.0057 239,745
Table 5 presents the mean and standard deviation of the test area under the receiver operating characteristic curve (ROC-AUC) for SIR-GCN and other models in literature. The tricks used and the number of parameters are also presented for completeness. The results show that with only a single GNN layer, SIR-GCN outperforms comparable models in predicting molecules inhibiting HIV replication, highlighting its expressivity. Given the simplicity of SIR-GCN, it is expected to exhibit lower performance compared to complex models such as the graph stochastic attention (GSAT) (Miao et al., 2022), which builds upon PNA by leveraging the information bottleneck principle, and the cell isomorphism network (CIN) (Bodnar et al., 2021), which is a hierarchical message-passing framework utilizing the topological features of graphs. Despite its simple design, the expressivity of SIR-GCN is evident in its close performance to that of complex GNN frameworks.

### v4

@@ 4.2 GAT
$e_{u,v}=\boldsymbol{a_{\text{GAT}}^{\top}}~{}\textsc{LeakyReLU}\left(\boldsymbol{W_{Q,\text{GAT}}}~{}\boldsymbol{h_{u}}+\boldsymbol{W_{K,\text{GAT}}}~{}\boldsymbol{h_{v}}\right),$ (16)
with the message from node $v$ to node $u$ proportional to $\exp\left(e_{u,v}\right)\cdot\boldsymbol{W_{K,\text{GAT}}}~{}\boldsymbol{h_{v}}$. While the attention mechanism of GATv2 is anisotropic and dynamic, its messages are nevertheless only linearly transformed with the query node $u$ only determining the degree of contribution of each message through the scalar $e_{u,v}$. Meanwhile, SIR-GCN directly applies the anisotropic and dynamic function in Eqn. 16 to the message function, allowing the features of the query node to dynamically transform messages. Specifically, if $\boldsymbol{W_{Q}}=\boldsymbol{W_{Q,\text{GAT}}}$, $\boldsymbol{W_{K}}=\boldsymbol{W_{K,\text{GAT}}}$, $\sigma=\textsc{LeakyReLU}$ and $\boldsymbol{W_{R}}=\boldsymbol{a_{\text{GAT}}^{\top}}$, one obtains
$\boldsymbol{h_{u}^{*}}=\sum_{v\in\mathcal{N}(u)}\boldsymbol{a_{\text{GAT}}^{\top}}~{}\textsc{LeakyReLU}\left(\boldsymbol{W_{Q,\text{GAT}}}~{}\boldsymbol{h_{u}}+\boldsymbol{W_{K,\text{GAT}}}~{}\boldsymbol{h_{v}}\right)$ (17)
which shows Eqn. 16 becoming a contextualized message in SIR-GCN. Nevertheless, GAT and GATv2 may also be recovered, up to a normalizing constant, with the appropriate parameters.
@@ 4.3 GIN
Likewise, within the proposed SIR-GCN, one may explicitly add a residual connection in the combination strategy to obtain
$\boldsymbol{h_{u}^{*}}=\textsc{MLP}_{\text{Res}}(\boldsymbol{h_{u}})+\sum_{v\in\mathcal{N}(u)}\boldsymbol{W_{R}}~{}\sigma\left(\boldsymbol{W_{Q}}\boldsymbol{h_{u}}+\boldsymbol{W_{K}}\boldsymbol{h_{v}}\right),$ (18)
where $\textsc{MLP}_{\text{Res}}$ is a learnable residual network. If $\textsc{MLP}_{\text{Res}}(\boldsymbol{h})=(1+\epsilon)\cdot\boldsymbol{h}$, $\sigma=\textsc{PReLU}(\alpha=1)$, $\boldsymbol{W_{Q}}=\boldsymbol{0}$, and $\boldsymbol{W_{R}W_{K}}=\boldsymbol{I}$, then
$\boldsymbol{h^{*}_{u}}=\left(1+\epsilon\right)\cdot\boldsymbol{h_{u}}+\sum_{v\in\mathcal{N}(u)}\boldsymbol{h_{v}}$ (19)
is equivalent to GIN. Hence, SIR-GCN with residual connection generalizes GIN.
@@ 4.4 PNA
Furthermore, while SIR-GCN and PNA approach the problem of uncountable node features differently, both models highlight the significance of anisotropic message functions considering both the features of the query and key nodes. The key difference lies with PNA using multiple aggregators and a static (Brody et al., 2022) (i.e., a function approximator with limited expressivity; e.g., linear) message function
$m\left(\boldsymbol{h_{v}},\boldsymbol{h_{u}}\right)=\boldsymbol{W_{K}}\boldsymbol{h_{v}}+\boldsymbol{W_{Q}}\boldsymbol{h_{u}}\eqcolon\boldsymbol{W_{K}}\boldsymbol{h_{v}}+\boldsymbol{b_{u}}.$ (20)
Consequently, the influence of the query node on the aggregated neighborhood feature becomes limited. In particular, when using mean, max, or min aggregators, the influence of the query node $u$ is restricted to the bias term $\boldsymbol{b_{u}}\coloneq\boldsymbol{W_{Q}}\boldsymbol{h_{u}}$. Moreover, with normalized moment aggregators, the bias term is effectively canceled out during the normalization process, further reducing the influence of the query node. Hence, PNA does not fully leverage its anisotropic nature, attributed to its heuristic application of multiple aggregators and scalers in a linear MPNN, thereby limiting its expressivity. In contrast, the dynamic nature of SIR-GCN allows for the non-linear and contextualized embedding of the features of the query node within the messages, thereby fully leveraging its anisotropic nature while allowing it to employ only a single aggregator.
@@
where the modification lies in concatenating the label of the query node $u$ with every element of the multiset before hashing. This modification, while negligible when $\mathcal{H}$ is countable, becomes significant when $\mathcal{H}$ is uncountable as highlighted in the previous section. Thus, SIR-GCN inherits the theoretical capabilities of the 1-WL test.
@@ 4.6 SIR-GCN
Overall, SIR-GCN is a simple MPNN instance that offers flexibility in two key dimensions of GNN—message transformation and aggregator. Consequently, it generalizes four classical GNNs in literature—GCN, GraphSAGE, GAT, and GIN—ensuring that it is at least as expressive as these models. Notably, SIR-GCN distinguishes itself from other GNNs by employing both anisotropic and dynamic (i.e., contextualized) messages within the MPNN framework, enabling the non-uniform aggregation of neighboring nodes in heterophilous graphs (Zheng et al., 2024) while maintaining adaptability to homophilous graphs.
@@
Figure 3: SIR-GCN generalizes classical GNNs.
In addition, SIR-GCN also distinguishes itself from PNA in addressing the problem of uncountable node features by employing only a single aggregator that theoretically holds for graphs of arbitrary sizes, thus reducing computational complexity. Nevertheless, its expressivity is maintained through contextualized messages via the application of only an activation function along edges, allowing it to inherit the representational capability of the 1-WL test.
@@ 5 Experiments
Experiments on synthetic and benchmark datasets in node and graph property prediction tasks are performed to highlight the expressivity of SIR-GCN. Following the evaluation methodology of Xu et al. (2019), Corso et al. (2020), and Brody et al. (2022), the key GNNs in literature without advanced architectural design nor domain-specific features are used as primary comparisons to ensure a fair evaluation.
@@ 5.1 Synthetic datasets
DictionaryLookup (Brody et al., 2022) consists of bipartite graphs with $2n$ $n$ key nodes each with an attribute and value and $n$ query nodes each with an attribute. The task is to predict the values of query nodes by matching their attributes with the key nodes as seen in Fig. 4.
@@
Table 1 presents the mean and standard deviation of the test accuracy for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $n$. Notably, SIR-GCN and GATv2 can achieve perfect accuracy in this synthetic task attributed to their anisotropic and dynamic nature, enabling them to learn the relationship between every query and key node. Nonetheless, it may be observed that GATv2 suffers from performance degradation in some trials. Meanwhile, the other models fail to predict the value of query nodes even for the training graphs due to their isotropic and/or static nature, hindering their ability to learn relationships between neighboring nodes. The results hence underscore the utility of a dynamic attentional or relational mechanism in capturing the relationship between the query and key nodes.
HeteroEdgeCount
HeteroEdgeCount is an original synthetic dataset consisting of randomly generated directed graphs with each node randomly labeled one of $c$ classes. The task is then to count the number of heterophilous directed edges in each graph connecting nodes with different class labels as illustrated in Fig. 5. This dataset is explicitly designed to highlight the limitations of key GNNs in literature, particularly the theoretically grounded GNNs such as GIN and PNA, even in trivial tasks involving countable node features. Specifically, it underscores the utility of anisotropic and dynamic message functions in learning the relationships between neighboring nodes. Crucially, this dataset is not intended to assess the ability of GNNs to handle heterophilous graphs, as this falls beyond the scope of this work.
Table 2: Test mean squared error on HeteroEdgeCount.
@@
Table 2 presents the mean and standard deviation of the test mean squared error (MSE) for SIR-GCN, GCN, GraphSAGE, GATv2, GIN, and PNA across different values of $c$. Notably, SIR-GCN consistently achieves near-zero MSE loss due to its anisotropic and dynamic nature as well as its sum aggregation, allowing it to learn the relationship between the labels of neighboring nodes while retaining the graph structure. In fact, for $\boldsymbol{W_{Q}}=\boldsymbol{I}$, $\boldsymbol{W_{K}}=-\boldsymbol{I}$, $\sigma=\textsc{ReLU}$, and $\boldsymbol{W_{R}}=\boldsymbol{1}^{\top}$, it may be shown that SIR-GCN will always produce the correct output for every graph. In contrast, GCN, GraphSAGE, and GATv2 obtained large MSE losses due to their mean or max aggregation which fails to preserve the graph structure as noted by Xu et al. (2019). Meanwhile, GIN and PNA successfully retain the graph structure with their sum aggregation but fail to differentiate neighboring nodes due to their static nature. The results thus illustrate the utility of anisotropic and dynamic message functions using sum aggregation even in simple tasks with countable node features, highlighting the limitations of existing GNNs.
@@ 5.2 Benchmark datasets
Benchmarking GNNs (Dwivedi et al., 2023) is a collection of benchmark datasets consisting of diverse mathematical and real-world graphs across various GNN tasks. In particular, the WikiCS, PATTERN, and CLUSTER datasets fall under node property prediction tasks while the MNIST, CIFAR10, and ZINC datasets fall under graph property prediction tasks. Furthermore, the WikiCS, MNIST, and CIFAR10 datasets have uncountable node features while the remaining datasets have countable node features. The performance metric of ZINC is the mean absolute error (MAE) while the performance metric of the remaining datasets is accuracy. These six benchmark datasets encompass a diverse range of GNN tasks, enabling a comprehensive and robust evaluation of model performance. Dwivedi et al. (2023) provides more information regarding the individual datasets.
@@
Note: Missing values indicate that no results were published in previous works.
Table 3 presents the mean and standard deviation of the test performance for SIR-GCN, GCN, GraphSAGE, GAT, GATv2, GIN, and PNA across the six benchmark datasets with the experimental set-up, such as parameter count and model architecture, following that of Dwivedi et al. (2023), Corso et al. (2020), and Tailor et al. (2022) to ensure a fair evaluation where performance differences are solely attributed to the GNN architectural design. The test performance for the efficient graph convolution single (EGC-S) and efficient graph convolution multiple (EGC-M) (Tailor et al., 2022) are also presented as additional MPNN-based baselines. Notably, SIR-GCN consistently outperforms classical GNNs in literature by a substantial margin which may be attributed to its ability to generalize these models, complementing the mathematical discussions in the previous section. Moreover, despite employing multiple aggregators and incurring higher computational complexity to ensure injectivity, PNA still fails to outperform the simpler and more computationally efficient SIR-GCN on datasets with uncountable node features, even though the former is explicitly designed for such tasks. Furthermore, SIR-GCN also outperforms the more recent EGC-S and EGC-M despite their use of additional tricks, including multiple convolutional basis weights, regularization heads, and aggregators. Overall, the results underscore that, under the same constraints, SIR-GCN consistently outperforms MPNN-based baselines despite its simplicity, establishing it as a promising alternative to existing MPNNs.
Open Graph Benchmark
Open Graph Benchmark (Hu et al., 2020a) is another collection of datasets consisting of realistic, large-scale, and diverse benchmarks for GNNs. In particular, the ogbn-arxiv with its uncountable node features falls under node property prediction tasks. Meanwhile, the ogbg-molhiv with its countable node features falls under graph property prediction tasks. The performance metric of ogbn-arxiv is accuracy while the performance metric of ogbg-molhiv is the area under the receiver operating characteristic curve (ROC-AUC). Hu et al. (2020a) provides more information regarding the individual datasets.
Table 4: Test performance on Open Graph Benchmark.
Model ogbn-arxiv ($\uparrow$) ogbg-molhiv ($\uparrow$)
GCN 71.92 $\pm$ 0.21 76.14 $\pm$ 1.29
GraphSAGE 71.73 $\pm$ 0.26 75.97 $\pm$ 1.69
GAT 71.81 $\pm$ 0.23 77.17 $\pm$ 1.37
GATv2 71.87 $\pm$ 0.43 77.15 $\pm$ 1.55
GIN 67.33 $\pm$ 1.47 76.02 $\pm$ 1.35
PNA 71.21 $\pm$ 0.30 79.05 $\pm$ 1.32
EGC-S 72.21 $\pm$ 0.17 77.44 $\pm$ 1.08
EGC-M 71.96 $\pm$ 0.23 78.18 $\pm$ 1.53
SIR-GCN 72.52 $\pm$ 0.16 77.63 $\pm$ 0.84
Table 4 presents the mean and standard deviation of the test performance for SIR-GCN, GCN, GraphSAGE, GAT, GATv2, GIN, PNA, EGC-S, and EGC-M across the two large-scale benchmark datasets with the experimental set-up following that of Corso et al. (2020) and Tailor et al. (2022) to ensure a fair evaluation. Notably, SIR-GCN still outperforms both classical GNNs and EGC-S by a significant margin even in these large-scale graphs, which may be attributed to its anisotropic and dynamic message function. Unsurprisingly, however, PNA and EGC-M exhibit better performance relative to SIR-GCN on ogbg-molhiv as this molecular property prediction task greatly benefits from maintaining graph isomorphism via multiple aggregators, scalers, convolutional basis weights, and regularization heads. Nevertheless, SIR-GCN outperforms both PNA and EGC-M by a substantial margin on ogbn-arxiv as this 40-class node classification task does not require maintaining graph isomorphism. Overall, the results highlight the significance of contextualized messages in enhancing GNN expressivity and the utility of softly relaxing the injective and metric requirements within the MPNN framework for most practical GNN applications.